# Interstitials as a key ingredient for P segregation to grain boundaries in polycrystalline $\alpha$-Fe

### v1

@@ Abstract
The segregation of solutes to grain boundaries can significantly influence material behavior. Most previous computational studies have concentrated on substitutional solute segregation, neglecting interstitial segregation due to its increased complexity. The site preference, interstitial or substitutional, for P segregation in $\alpha$-Fe still remains under debate. In this work, we investigate the full GB-segregation spectrum for both substitutional and interstitial GB sites in a polycrystalline atomistic structure of ferrite with the aid of classical interatomic potentials combined with machine learning techniques. The method is qualitatively tested for H and Ni, where the segregation behavior in $\alpha$-Fe is well understood. Our findings for P show that segregation to both types of GB sites is possible, with a preference for the substitutional sites based on the mean segregation energy. However, due to the much larger number of interstitial sites, interstitial segregation significantly contributes to the GB enrichment with P. This underscores the importance of considering interstitial P segregation in addition to the substitutional one. Furthermore, we also argue that equally important for quantitative predictions (that agree with experimental data) is to get a representative spectrum of the segregation energies.
@@
journal: Acta Materialia
@@ 1 Introduction
In this work, we aim to resolve the ongoing debate regarding the site preference of P at GBs in $\alpha$-Fe by investigating the segregation spectrum for both interstitial and substitutional P. First, we seek to clarify the site preference of P at GBs. Second, we make use of the complete segregation profile—including both interstitial and substitutional segregation energies—to calculate GB concentrations at different temperatures and bulk concentrations. These results are then compared with the experimental findings of Erhart and Grabke [13] and the computational results of Reichmann et al. [2]. To achieve this, we utilize recent advances in machine learning techniques for predicting segregation energy spectra [14] and extend these methods to predict interstitial segregation energies in body-centered cubic materials. We also predict the segregation profiles for H and Ni. For these solutes, the segregation behavior is well-established in the literature, providing the possibility to validate our approach.
2 Methodology
2.1 Atomistic model of polycrystal
@@ 2.2 Segregation energy spectra
We note that the segregation energies were calculated only for transitions from the most stable bulk site, i.e., tetrahedral for H and substitutional for P and Ni, to the GB site of interest, which could be either substitutional or interstitial. Hence, the calculation of the segregation energy $\Delta E_{seg}^{k}(X)$ of solute $X$ to site $k$ can be summarized as follows:
@@ 2.3 Thermodynamic modeling of GB enrichment
where the fraction of GB sites is given by the factor $f_{GB}$, which can reasonably be assumed to be $610^{-}$ 6 for large grain diameters [2]. To obtain an average GB solute concentration $\bar{C}_{GB}(X)$ of solute $X$, the total amount of segregated P atoms is calculated from a summation over $C_{GB}^{k}(X)$ of all sites (interstitial and substitutional combined) and divided by the total number of atoms in the GB (Fe and P) $N_{GB}$[5]:
@@
3 Results and discussion
3.1 Interstitial and substitutional segregation profiles
The performance of the machine learning models in predicting both substitutional and interstitial segregation energies is shown in Fig. 2. A clear trend in accuracy can be observed. For substitutional solute segregation in Fe, Ni exhibits the highest accuracy, followed by P and H, whereas for interstitial solute segregation, the order is reversed. This suggests that the greater the size mismatch between the GB site size and the atomic radius of the solute (i.e., the more structural relaxation taking place), the higher the uncertainty in the model’s predictions.
@@ 3.1 Interstitial and substitutional segregation profiles
Figure 2: Performance of the ML model for segregation energy predictions. The predicted segregation energies are plotted against the calculated segregation energies for P (left column), Ni (middle), and H (right) in the substitutional (top row) and interstitial (bottom row) GB sites. The dashed line indicates a perfect fit.
@@
It is evident that the segregation behavior of Ni is markedly distinct. The substitutional segregation energies exhibit a mean value of approximately $-0.1\,\text{eV}$, indicating a somewhat weaker driving force for segregation than what is found for P. Notably, all interstitial sites demonstrate anti-segregating behavior with positive segregation energies. It can thus be concluded that Ni atoms will segregate exclusively to substitutional sites. The segregation energy distribution of H exhibits a contrasting behavior. The segregation energy to mean interstitial sites is $-0.16\,\text{eV}$, whereas all substitutional segregation energies are positive.
Table 2: Summary of DFT calculated segregation energies from literature for P, Ni, and H in different special coincident site lattice GBs (rows). Parentheses indicate segregation energies given as a function of GB distance. The superscript letter S denotes substitutional segregation, while the superscript letter I denotes interstitial segregation.
GB P Ni H
$\Sigma 3(1\bar{1}1)[110]$ (0.10(0.10( 0.10, $0.04$, $0.00$, $-0.02$, $-0.04$, $-0.05$, $-0.10$, 0.33)-0.33)- 0.33 )S[8] (0.00(0.00( 0.00, $-0.08$, $-0.09$, $-0.18$, 0.39)-0.39)- 0.39 )S[6] (0.36(-0.36( - 0.36, 0.48)-0.48)- 0.48 )I[28]
$\Sigma 3(1\bar{1}2)[110]$ (0.03(0.03( 0.03, $-0.01$, $-0.01$, $-0.01$, $-0.17$, $-0.18$, 0.19)-0.19)- 0.19 )S[8] (0.02(-0.02( - 0.02, $-0.03$, $-0.12$, $-0.13$, 0.33)-0.33)- 0.33 )S[6] --- -
$\Sigma 5(012)[100]$ (0.16(0.16( 0.16, $-0.05$, $-0.54$ $-0.71$, $-0.92$S[5], (0.597(-0.597( - 0.597, $-0.635$, 0.604)-0.604)- 0.604 )I[5] --- - (0.34(-0.34( - 0.34, $-0.37$, 0.42)-0.42)- 0.42 )I[28]
$\Sigma 9(2\bar{2}1)[110]$ (0.24(0.24( 0.24, $0.01$, $-0.11$, $-0.15$, $-0.17$, $-0.53$, $-0.84$, $-0.95$, 1.37)-1.37)- 1.37 )S[8] $0.06$, $-0.01$, $-0.10$, $-0.10$, $-0.11$, $-0.12$ $-0.12$, $-0.40$, $-0.43$S[6] --- -
$\Sigma 11(3\bar{3}2)[110]$ (0.04(0.04( 0.04, $0.02$, $0.01$, $-0.01$, $-0.05$, $-0.06$, $-0.07$, $-0.08$, $-1.03$, $-1.03$, $-1.03$, 1.05)-1.05)- 1.05 )S[8] (0.00(0.00( 0.00, $-0.03$, $-0.06$, $-0.07$, $-0.20$, $-0.20$, $-0.22$, $-0.22$, $-0.56$, $-0.56$, 0.56)-0.56)- 0.56 )S[8]
$\Sigma 13(015)[100]$ (0.40(0.40( 0.40, $0.03$, $-0.08$, $-0.35$, $-0.37$, 1.01)-1.01)- 1.01 )S[5], $-1.35$I[5]
$\Sigma 17(014)[100]$ (0.38(0.38( 0.38, $0.07$, $-0.00$, $-0.08$, $-0.28$, $-0.29$, $-0.34$, $-0.83$, 1.00)-1.00)- 1.00 )S[5], $-1.45$, $-0.8$I[5]
A comprehensive overview of literature data of DFT calculated segregation energies to different coincidence site lattice GBs for these solutes is given in Table 2. From this, it can be seen that a general consensus in existing literature exists that classical substitutional elements in bulk, such as Ni, Cr, Mo, etc, also segregate to substitutional sites at GBs. The same consensus exists for small interstitial atoms such as C and H and their segregation to interstitial sites. These trends are perfectly reproduced by our results showing that H does not segregate to substitutional sites and Ni does not segregate to interstitial sites. Further, the absolute values given in Table 2 agree well with our results for segregation in polycrystalline Fe. This indicates the reliability of the here proposed approach using Voronoi tesselation with MD and ML for obtaining the full interstitial segregation spectrum.
@@
3.2 GB enrichment of P
With the P segregation spectra in hand, we proceed with the evaluation of the average GB solute concentrations ($\bar{C}_{GB}$) using Eq. 4 and compare these with the experimental findings of Erhart and Grabke [13]. In general, the calculated GB concentrations increase with increasing P content and decrease with increasing temperature, in agreement with the experiments, as can be seen in Fig. 4. However, the quantitative agreement with the experimental measurements clearly depends on the segregation spectra taken into account.
@@ 3.2 GB enrichment of P
Note that interstitial sites modify the average GB concentration not only through the different segregation spectrum but also via the denominator in equation 4. For Fe-P, this has a general tendency to flatten the curve of $\bar{C}_{GB}(P)$ as a function of $T$. Together with the complete segregation spectra (both substitutional and interstitial), it brings the computational result into good agreement with the experimental data.
@@
4 Conclusions
In this work, we analyzed the substitutional and interstitial segregation spectra of P in $\alpha$-Fe using a polycrystalline GB model. Interstitial sites were selected based on Voronoi polyhedral construction, resulting in approximately $1.19\times 10^{6}$ potential segregation sites in the GB regions, composed of $1.08\times 10^{5}$ substitutions (lattice) sites. The complete segregation energy spectrum was achieved with the help of a machine learning model based on explicit calculations of segregation energies of 100/300 representative substitutional/interstitial segregation sites using the classical interatomic potential. For comparison, we also examined the segregation behavior of Ni and H, elements for which the existing literature predominantly assumes segregation to substitutional and interstitial sites, respectively. The main findings of this study are as follows:
@@ 4 Conclusions
5 Acknowledgment

### v2

@@ Abstract
Solute segregation to grain boundaries (GBs) significantly impacts material behavior, with most studies focusing on substitutional solute segregation while neglecting interstitial segregation due to its increased complexity. The site preference, interstitial or substitutional, for P segregation in $\alpha$-Fe still remains under debate. This work investigates both substitutional and interstitial GB segregation in a polycrystalline model using classical interatomic potentials and machine learning. The method is validated with H and Ni, whose segregation behaviors are well understood. For P, we find segregation to both GB site types, with a preference for substitutional sites based on mean segregation energy. However, the abundance of interstitial sites means interstitial segregation also significantly contributes to the GB enrichment with P. This highlights the importance of considering interstitial P segregation alongside substitutional segregation. Additionally, obtaining a representative spectrum of segregation energies is crucial for accurate, experimentally aligned predictions.
@@
journal: Scripta Materialia
@@
In this work, we aim to resolve the ongoing debate regarding the site preference of P at GBs in $\alpha$-Fe by investigating the segregation spectrum for both interstitial and substitutional P in a polycrystalline model. First, we seek to clarify the site preference of P at GBs. Second, we make use of the complete segregation profile—including both interstitial and substitutional segregation energies—to calculate GB concentrations at different temperatures and bulk concentrations. These results are then compared with the experimental findings of Erhart and Grabke [13] and the computational results of Reichmann et al. [2]. To achieve this, we utilize recent advances in machine learning techniques for predicting segregation energy spectra [14] and extend these methods to predict interstitial segregation energies in body-centered cubic materials. We also predict the segregation profiles for H and Ni. For these solutes, the segregating behavior is well established in the literature, providing an opportunity to validate the approach.
@@
We note that the segregation energies were calculated only for transitions from the most stable bulk site (see Supplementary Material Fig. S1 for the overview of solute formation energies), i.e., tetrahedral for H and substitutional for P and Ni, to the GB site of interest, which could be either substitutional or interstitial. Hence, the calculation of the segregation energy $\Delta E_{seg}^{k}(X)$ of solute $X$ to site $k$ can be summarized as follows:
@@
where the fraction of GB sites is given by the factor $f_{GB}$, which can reasonably be assumed to be $10^{-6}$ for large grain diameters [2]. To obtain an average GB solute concentration $\bar{C}_{GB}(X)$ of solute $X$, the total amount of segregated P atoms is calculated from a summation over $C_{GB}^{k}(X)$ of all sites (interstitial and substitutional combined) and divided by the total number of atoms in the GB (Fe and P) $N_{GB}$[5]:
@@
The performance of the machine learning models in predicting both substitutional and interstitial segregation energies, based on the training data set ($\equiv$ clustering points) is shown in Fig. 2. A clear trend in accuracy can be observed. For substitutional solute segregation in Fe, Ni exhibits the highest accuracy, followed by P and H, whereas for interstitial solute segregation, the order is reversed. This suggests that the greater the size mismatch between the GB site size and the atomic radius of the solute (i.e., the more structural relaxation taking place), the higher the uncertainty in the model’s predictions.
@@
Figure 2: Performance of the ML model for segregation energy predictions. The shown points and the respective RMSE values are based on the training data set. The predicted segregation energies are plotted against the calculated segregation energies for P (left column), Ni (middle), and H (right) in the substitutional (top row) and interstitial (bottom row) GB sites. The dashed line indicates a perfect fit.
@@
It is evident that the segregation behavior of Ni is markedly distinct. The substitutional segregation energies exhibit a mean value of approximately $-0.1\,\text{eV}$, indicating a somewhat weaker driving force for segregation than what is found for P. Notably, all interstitial sites demonstrate anti-segregating behavior with positive segregation energies. It can thus be concluded that Ni atoms will segregate exclusively to substitutional sites. The segregation energy distribution of H exhibits a contrasting behavior. The segregation energy to mean interstitial sites is $-0.16\,\text{eV}$, whereas all substitutional segregation energies are positive. Another difference emerges when comparing the H bulk formation energies. Supplementary Material Fig. S1 shows that the tetrahedral position is only slightly more stable ($\approx 0.05\,\text{eV}$ for the used EAM potential) than the octahedral interstitial position. Since a typical unit of thermal energy at room temperature is $k_{B}T\approx 0.026\,\text{eV}$, these results suggest that octahedral bulk positions can serve as potential starting points for H segregation. The cumulative segregation spectrum is shown in Supplementary Material Fig. S4. However, the change is small, shifting the mean segregation energy from $-0.16\,\text{eV}$ to $-0.17\,\text{eV}$. Also noteworthy are the multiple peaks in the distribution of the interstitial segregation spectra of P and H. Similar work investigating the segregation behavior of H in fcc Ni [28] and fcc Pd [29], respectively, concluded that these belong to different groups of interstitial sites with characteristic fingerprints divided into sites located in the core and at the surface of the GB, respectively.
A comprehensive overview of literature data of DFT calculated segregation energies to different coincidence site lattice GBs for these solutes is provided in graphical and tabular form in Supplementary Material Fig. S2 and Table S1. From this, it can be seen that a general consensus in existing literature exists that classical substitutional elements in bulk, such as Ni, Cr, Mo, etc, also segregate to substitutional sites at GBs. The same consensus exists for small interstitial atoms such as C and H and their segregation to interstitial sites. These trends are perfectly reproduced by our results showing that H does not segregate to substitutional sites and Ni does not segregate to interstitial sites. This indicates the reliability of the here proposed approach using Voronoi tesselation with MD and ML for obtaining the full interstitial segregation spectrum.
@@
With the P segregation spectra in hand, we proceed with the evaluation of the average GB solute concentrations ($\bar{C}_{GB}$) using Eq. 4 and compare these with the experimental findings of Erhart and Grabke [13]. Analogous segregation isotherms for Fe-Ni and Fe-H systems can be found in Supplementary Material, Fig. S5. In general, the calculated GB concentrations increase with increasing P content and decrease with increasing temperature, in agreement with the experiments, as can be seen in Fig. 4. However, the quantitative agreement with the experimental measurements clearly depends on the segregation spectra taken into account.
@@
Note that interstitial sites modify the average GB concentration not only through the different segregation spectrum but also via the denominator in Eq. 4. For Fe-P, this has a general tendency to flatten the curve of $\bar{C}_{GB}(P)$ as a function of $T$. Together with the complete segregation spectra (both substitutional and interstitial), it brings the computational result into good agreement with the experimental data. Finally, solute-solute interactions between the different site types may influence the agreement with experiments, particularly at higher solute concentrations. However, due to computational complexity, these interactions are not considered here. Investigating their potential impact on the segregation isotherms is left to future work.
@@
In conclusion, we analyzed the substitutional and interstitial segregation spectra of P in $\alpha$-Fe using a polycrystalline GB model. Interstitial sites were selected based on Voronoi polyhedral construction, resulting in approximately $1.19\times 10^{6}$ potential segregation sites in the GB regions, in addition to $1.08\times 10^{5}$ substitutional (lattice) sites. The complete segregation energy spectrum was achieved with the help of a machine learning model based on explicit calculations of segregation energies of 100/300 representative substitutional/interstitial segregation sites using classical interatomic potentials. For comparison, we also examined the segregation behavior of Ni and H, elements for which the existing literature predominantly assumes segregation to substitutional and interstitial sites, respectively. The main findings of this study are as follows:
@@
  • 4.
    It is found that the accuracy of the applied machine learning method using SOAP features based on the undecorated GB model depends strongly on the amount of structural relaxation induced by the segregating species. Thus, the accuracy for substitutional solute segregation decreases in the order of Ni, P, H and for interstitial solute segregation in the order of H, P, Ni, represented by the size mismatch between solute and GB site.
Acknowledgment